OpenStudy (anonymous):

Part 1: Explain, using complete sentences, how to simplify the expression below. Part 2: What is the simplified expression? (3x3y4)(2x2y6)

OpenStudy (mayankdevnani):

\[3\times 3y^4=9y^4\]

OpenStudy (mayankdevnani):

\[2\times 2y^6=4y^6\]

OpenStudy (mayankdevnani):

\[(9y^4)\times (4y^6)=36y ^10\]

OpenStudy (mayankdevnani):



OpenStudy (mayankdevnani):

ok @debbies

OpenStudy (anonymous):

my mistake is your equation like this?  if so then just do this do 3x2 which is 6 then (3y^4)(2y^10) = 6y^10 then multiply those 2 together and get 36y^10